# BRIGHT: A globally distributed multimodal building damage assessment dataset with very-high-resolution for all-weather disaster response

### v1

@@ 4.3 Significance of BRIGHT
Hongruixuan Chen: conceptualization (lead), data curation (lead), funding acquisition (equal), methodology (lead), project administration (lead), investigation (lead), software (lead), writing – original draft preparation (lead), visualization (lead). Jian Song: conceptualization (support), data curation (equal), funding acquisition (support), investigation (support), writing – original draft preparation (equal), visualization (support). Olivier Dietrich: conceptualization (support), data curation (equal), software (support), visualization (equal), writing – review and editing (support). Clifford Broni-Bediako: conceptualization (support), data curation (support), writing – review and editing (equal). Weihao Xuan: data curation (support), project administration (support), writing – review and editing (support). Junjue Wang: data curation (support), writing – review and editing (support). Xinlei Shao: funding acquisition (support), writing – review and editing (support). Yimin Wei: writing – review and editing (support). Junshi Xia: conceptualization (support), data curation (support), supervision (support), writing – review and editing (support). Cuiling Lan444Cuiling Lan did not participate in any activities related to the acquisition, processing, utilization, or distribution of the datasets. Her access to the data was solely limited to the information presented in the current paper, similar to that of the readers. : conceptualization (support), supervision (support), writing – linguistic refinement (equal). Konrad Schindler: conceptualization (equal), resources (support), supervision (equal), writing – review and editing (equal). Naoto Yokoya: conceptualization (equal), funding acquisition (lead), resources (lead), supervision (lead), writing – review and editing (equal).
@@ Acknowledgements.
This work was supported in part by the JSPS, KAKENHI under Grant Number 24KJ0652 and 22H03609, the Council for Science, Technology and Innovation (CSTI), the Cross-ministerial Strategic Innovation Promotion Program (SIP), Development of a Resilient Smart Network System against Natural Disasters (Funding agency: NIED), JST, FOREST under Grant Number JPMJFR206S, Microsoft Research Asia, Next Generation AI Research Center of The University of Tokyo, and Young Researchers Exchange Programme between Japan and Switzerland under the Japanese-Swiss Science and Technology Programme. The authors would also like to give special thanks to Sarah Preston of Capella Space, Capella Space’s Open Data Gallery, Maxar Open Data Program and Umbra Space’s Open Data Program for providing the valuable data.References
@@
- Chen et al. (2022) Chen, H., Nemni, E., Vallecorsa, S., Li, X., Wu, C., and Bromley, L.: Dual-Tasks Siamese Transformer Framework for Building Damage Assessment, in: International Geoscience and Remote Sensing Symposium (IGARSS), pp. 1600–1603, 2022.
@@
- Sun et al. (2024) Sun, Y., Wang, Y., and Eineder, M.: QuickQuakeBuildings: Post-Earthquake SAR-Optical Dataset for Quick Damaged-Building Detection, IEEE Geoscience and Remote Sensing Letters, 21, 1–5, 2024.

### v2

@@ Appendix G Introduction of UMCD methods and a more practical evaluation protocol
Hongruixuan Chen: conceptualization (lead), data curation (lead), funding acquisition (equal), methodology (lead), project administration (lead), investigation (lead), software (lead), visualization (lead), writing – original draft preparation (lead), writing – review and editing (lead). Jian Song: conceptualization (support), data curation (equal), funding acquisition (support), methodology (support), investigation (support), software (support), writing – original draft preparation (equal), visualization (support). Olivier Dietrich: conceptualization (support), data curation (equal), software (support), visualization (equal), writing – review and editing (support). Clifford Broni-Bediako: conceptualization (support), data curation (support), writing – review and editing (equal). Weihao Xuan: data curation (support), project administration (support), writing – review and editing (equal). Junjue Wang: data curation (support), writing – review and editing (support). Xinlei Shao: funding acquisition (support), writing – review and editing (support). Yimin Wei: writing – review and editing (support). Junshi Xia: conceptualization (support), data curation (support), supervision (support), writing – review and editing (support). Cuiling Lan999Cuiling Lan did not participate in any activities related to the acquisition, processing, utilization, or distribution of the datasets. Her access to the data was solely limited to the information presented in the current paper, similar to that of the readers. : conceptualization (support), supervision (support), writing – linguistic refinement (equal). Konrad Schindler: conceptualization (equal), resources (support), supervision (equal), writing – review and editing (equal). Naoto Yokoya: conceptualization (equal), funding acquisition (lead), resources (lead), supervision (lead), writing – review and editing (equal).
@@ Acknowledgements.
This work was supported in part by the JSPS, KAKENHI under Grant Number 24KJ0652 and 22H03609, the Council for Science, Technology and Innovation (CSTI), the Cross-ministerial Strategic Innovation Promotion Program (SIP), Development of a Resilient Smart Network System against Natural Disasters (Funding agency: NIED), JST, FOREST under Grant Number JPMJFR206S, Microsoft Research Asia, Next Generation AI Research Center of The University of Tokyo, and Young Researchers Exchange Programme between Japan and Switzerland under the Japanese-Swiss Science and Technology Programme. The authors would also like to give special thanks to Sarah Preston of Capella Space, Capella Space’s Open Data Gallery, Maxar Open Data Program, and Umbra Space’s Open Data Program for providing the valuable data.References
@@
- Bastos Moroz and Thieken (2024) Bastos Moroz, C. and Thieken, A. H.: Urban growth and spatial segregation increase disaster risk: lessons learned from the 2023 disaster on the North Coast of São Paulo, Brazil, Natural Hazards and Earth System Sciences, 24, 3299–3314, 2024.
@@
- Chen et al. (2022a) Chen, H., Nemni, E., Vallecorsa, S., Li, X., Wu, C., and Bromley, L.: Dual-Tasks Siamese Transformer Framework for Building Damage Assessment, in: International Geoscience and Remote Sensing Symposium (IGARSS), pp. 1600–1603, 2022a.
- Chen et al. (2022b) Chen, H., Yokoya, N., Wu, C., and Du, B.: Unsupervised Multimodal Change Detection Based on Structural Relationship Graph Representation Learning, IEEE Trans. Geosci. Remote Sens., pp. 1–18, 2022b.
- Chen et al. (2023) Chen, H., Yokoya, N., and Chini, M.: Fourier domain structural relationship analysis for unsupervised multimodal change detection, ISPRS J. Photogramm. Remote Sens., 198, 99–114, 2023.
@@
- Chen et al. (2021) Chen, X., Yuan, Y., Zeng, G., and Wang, J.: Semi-Supervised Semantic Segmentation with Cross Pseudo Supervision, in: IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pp. 2613–2622, 2021.
@@
- Han et al. (2024) Han, T., Tang, Y., Zou, B., and Feng, H.: Unsupervised multimodal change detection based on adaptive optimization of structured graph, International Journal of Applied Earth Observation and Geoinformation, 126, 103 630, 2024.
- He et al. (2016) He, K., Zhang, X., Ren, S., and Sun, J.: Deep residual learning for image recognition, Proceedings of the IEEE Computer Society Conference on Computer Vision and Pattern Recognition (CVPR), pp. 770–778, 2016.
@@
- Ke et al. (2020) Ke, Z., Qiu, D., Li, K., Yan, Q., and Lau, R. W. H.: Guided Collaborative Training for Pixel-Wise Semi-Supervised Learning, in: European Conference Computer Vision, p. 429–445, 2020.
- Kreibich et al. (2022) Kreibich, H., Loon, A. F. V., Schröter, K., Ward, P. J., Mazzoleni, M., Sairam, N., Abeshu, G. W., Agafonova, S., AghaKouchak, A., Aksoy, H., Alvarez-Garreton, C., Aznar, B., Balkhi, L., Barendrecht, M. H., Biancamaria, S., Bos-Burgering, L., Bradley, C., Budiyono, Y., Buytaert, W., Capewell, L., Carlson, H., Cavus, Y., Couasnon, A., Coxon, G., Daliakopoulos, I., de Ruiter, M. C., Delus, C., Erfurt, M., Esposito, G., François, D., Frappart, F., Freer, J., Frolova, N., Gain, A. K., Grillakis, M., Grima, J. O., Guzmán, D. A., Huning, L. S., Ionita, M., Kharlamov, M., Khoi, D. N., Kieboom, N., Kireeva, M., Koutroulis, A., Lavado-Casimiro, W., Li, H.-Y., Llasat, M. C., Macdonald, D., Mård, J., Mathew-Richards, H., McKenzie, A., Mejia, A., Mendiondo, E. M., Mens, M., Mobini, S., Mohor, G. S., Nagavciuc, V., Ngo-Duc, T., Huynh, T. T. N., Nhi, P. T. T., Petrucci, O., Nguyen, H. Q., Quintana-Seguí, P., Razavi, S., Ridolfi, E., Riegel, J., Sadik, M. S., Savelli, E., Sazonov, A., Sharma, S., Sörensen, J., Souza, F. A. A., Stahl, K., Steinhausen, M., Stoelzle, M., Szalińska, W., Tang, Q., Tian, F., Tokarczyk, T., Tovar, C., Tran, T. V. T., Huijgevoort, M. H. J. V., van Vliet, M. T. H., Vorogushyn, S., Wagener, T., Wang, Y., Wendt, D. E., Wickham, E., Yang, L., Zambrano-Bigiarini, M., Blöschl, G., and Baldassarre, G. D.: The challenge of unprecedented floods and droughts in risk management, Nature, 608, 80–86, 2022.
@@
- Lian et al. (2019) Lian, Q., Duan, L., Lv, F., and Gong, B.: Constructing Self-Motivated Pyramid Curriculums for Cross-Domain Semantic Segmentation: A Non-Adversarial Approach, in: IEEE/CVF International Conference on Computer Vision (ICCV), pp. 6757–6766, 2019.
- Liu et al. (2025) Liu, T., Xu, J., Lei, T., Wang, Y., Du, X., Zhang, W., Lv, Z., and Gong, M.: AEKAN: Exploring Superpixel-Based AutoEncoder Kolmogorov-Arnold Network for Unsupervised Multimodal Change Detection, IEEE Transactions on Geoscience and Remote Sensing, 63, 1–14, 2025.
@@
- Liu et al. (2021) Liu, Z., Lin, Y., Cao, Y., Hu, H., Wei, Y., Zhang, Z., Lin, S., and Guo, B.: Swin Transformer: Hierarchical Vision Transformer using Shifted Windows, in: IEEE/CVF International Conference on Computer Vision (ICCV), pp. 9992–10 002, 2021.
@@
- Luo et al. (2019) Luo, Y., Zheng, L., Guan, T., Yu, J., and Yang, Y.: Taking a Closer Look at Domain Shift: Category-Level Adversaries for Semantics Consistent Domain Adaptation, in: IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pp. 2502–2511, 2019.
- Lv et al. (2022) Lv, Z., Huang, H., Li, X., Zhao, M., Benediktsson, J. A., Sun, W., and Falco, N.: Land Cover Change Detection With Heterogeneous Remote Sensing Images: Review, Progress, and Perspective, Proceedings of the IEEE, 110, 1976–1991, 2022.
@@
- Oquab et al. (2024) Oquab, M., Darcet, T., Moutakanni, T., Vo, H. V., Szafraniec, M., Khalidov, V., Fernandez, P., HAZIZA, D., Massa, F., El-Nouby, A., Assran, M., Ballas, N., Galuba, W., Howes, R., Huang, P.-Y., Li, S.-W., Misra, I., Rabbat, M., Sharma, V., Synnaeve, G., Xu, H., Jegou, H., Mairal, J., Labatut, P., Joulin, A., and Bojanowski, P.: DINOv2: Learning Robust Visual Features without Supervision, Transactions on Machine Learning Research, 2024.
- Ouali et al. (2020) Ouali, Y., Hudelot, C., and Tami, M.: Semi-Supervised Semantic Segmentation With Cross-Consistency Training, in: IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pp. 12 671–12 681, 2020.
@@
- Selvaraju et al. (2017) Selvaraju, R. R., Cogswell, M., Das, A., Vedantam, R., Parikh, D., and Batra, D.: Grad-CAM: Visual Explanations from Deep Networks via Gradient-Based Localization, in: 2017 IEEE International Conference on Computer Vision (ICCV), pp. 618–626, 2017.
@@
- Shi et al. (2015) Shi, X., Chen, Z., Wang, H., Yeung, D.-Y., Wong, W.-k., and Woo, W.-c.: Convolutional LSTM Network: a machine learning approach for precipitation nowcasting, in: Proceedings of the 29th International Conference on Neural Information Processing Systems, vol. 1, p. 802–810, 2015.
@@
- Sun et al. (2021) Sun, Y., Lei, L., Guan, D., and Kuang, G.: Iterative Robust Graph for Unsupervised Change Detection of Heterogeneous Remote Sensing Images, IEEE Trans. Image Process., 30, 6277–6291, 2021.
- Sun et al. (2024a) Sun, Y., Lei, L., Guan, D., Wu, J., Kuang, G., and Liu, L.: Image Regression With Structure Cycle Consistency for Heterogeneous Change Detection, IEEE Transactions on Neural Networks and Learning Systems, 35, 1613–1627, 2024a.
- Sun et al. (2024b) Sun, Y., Wang, Y., and Eineder, M.: QuickQuakeBuildings: Post-Earthquake SAR-Optical Dataset for Quick Damaged-Building Detection, IEEE Geoscience and Remote Sensing Letters, 21, 1–5, 2024b.
- Tarvainen and Valpola (2017) Tarvainen, A. and Valpola, H.: Mean teachers are better role models: Weight-averaged consistency targets improve semi-supervised deep learning results, in: Advances in Neural Information Processing Systems, edited by Guyon, I., Luxburg, U. V., Bengio, S., Wallach, H., Fergus, R., Vishwanathan, S., and Garnett, R., vol. 30, 2017.
@@
- Tsai et al. (2018) Tsai, Y.-H., Hung, W.-C., Schulter, S., Sohn, K., Yang, M.-H., and Chandraker, M.: Learning to Adapt Structured Output Space for Semantic Segmentation, in: 2018 IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 7472–7481, 2018.
@@
- van der Maaten and Hinton (2008) van der Maaten, L. and Hinton, G.: Visualizing Data using t-SNE, Journal of Machine Learning Research, 9, 2579–2605, 2008.
- Vu et al. (2019) Vu, T.-H., Jain, H., Bucher, M., Cord, M., and Pérez, P.: ADVENT: Adversarial Entropy Minimization for Domain Adaptation in Semantic Segmentation, in: IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pp. 2512–2521, 2019.
@@
- Xie et al. (2021) Xie, E., Wang, W., Yu, Z., Anandkumar, A., Alvarez, J. M., and Luo, P.: SegFormer: Simple and Efficient Design for Semantic Segmentation with Transformers, in: Advances in Neural Information Processing Systems (NIPS), vol. 34, pp. 12 077–12 090, 2021.
@@
- Yang et al. (2023) Yang, J., Qian, H., Xu, Y., Wang, K., and Xie, L.: Can We Evaluate Domain Adaptation Models Without Target-Domain Labels?, arXiv preprint arXiv:2305.18712, 2023.
- Yang and Soatto (2020) Yang, Y. and Soatto, S.: FDA: Fourier Domain Adaptation for Semantic Segmentation, in: IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pp. 4084–4094, 2020.
@@
- Ye et al. (2017) Ye, Y., Shen, L., Hao, M., Wang, J., and Xu, Z.: Robust Optical-to-SAR Image Matching Based on Shape Properties, IEEE Geoscience and Remote Sensing Letters, 14, 564–568, 2017.
- Ye et al. (2019) Ye, Y., Bruzzone, L., Shan, J., Bovolo, F., and Zhu, Q.: Fast and Robust Matching for Multimodal Remote Sensing Image Registration, IEEE Transactions on Geoscience and Remote Sensing, 57, 9059–9070, 2019.